# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ 13.B.3 Procedure understanding
Keystep assignment is applied to obtain a pseudo-labeling of the provided video segments when the supervision is at the procedure-level (i.e., segments are unlabeled and only keystep names are provided). This is achieved by means of an EgoVLPv2 model [85] pre-trained on ego-exo videos and narrations. Video segments and keystep names are projected to the shared video-language space using EgoVLPv2. We hence assigned each video segment to the closest keystep in the representation space according to the cosine distance. In the problem formulation with instance-level supervision (i.e., when keystep labels are available for all segments during both training and testing), we use ground truth labels instead of those obtained from keystep assignment.
- Uniform Baseline - Val (v1/v2) 58.99/59.12 58.11/57.04 59.07/58.02 65.05/67.37 65.47/65.71
Instance-Level Graph-Based Ground Truth Val (v1/v2) 82.11/82.11 58.96/58.97 72.11/72.23 84.12/84.12 63.91/63.92
Instance-Level End-to-End Ground Truth Val (v1/v2) 55.13/55.12 51.48/51.49 54.37/54.38 54.00/53.99 54.12/54.38
Procedure-Level Graph-Based Keystep Assignment Val (v1/v2) 72.66/70.47 70.19/68.67 89.17/80.49 73.34/70.85 60.01/58.56
Procedure-Level End-to-End Keystep Assignment Val (v1/v2) 54.20/54.18 53.55/53.56 53.95/53.96 52.25/52.23 55.44/55.42
 - Uniform Baseline - Test (v1/v2) 59.12/58.95 57.93/58.15 58.79/60.53 65.27/65.66 65.49/65.97
Instance-Level Graph-Based Ground Truth Test (v1/v2) 81.66/81.82 60.54/62.30 71.83/73.99 82.00/82.54 62.58/63.38
Instance-Level End-to-End Ground Truth Test (v1/v2) 51.85/51.29 70.31/66.01 52.01/51.99 51.73/51.75 50.94/51.23
Procedure-Level Graph-Based Keystep Assignment Test (v1/v2) 63.60/67.33 68.49/68.64 78.23/69.80 63.64/68.01 57.57/55.73
Procedure-Level End-to-End Keystep Assignment Test (v1/v2) 52.15/52.48 68.64/64.14 52.42/52.83 50.84/51.38 50.50/50.75
Table 17: Results for the procedure understanding task. All models are trained on V1. Best results are reported in bold, the second best results are underlined. All results are in percentage.
The procedural reasoning component creates for each procedure a transition graph based on keystep co-occurrences. In the graph, each node represents a keystep category, while directed edges represent the probability of transitioning from one node to another one. An edge $A\to B$ is assigned the following weight based on statistics collected from the training videos:
@@
End-to-end baseline. This baseline aims to provide an end-to-end approach to perform procedure understanding directly from the input clip. The baseline predicts previous keysteps, optional keysteps, and next keysteps by feeding video segment features extracted with EgoVLPv2 to three dedicated MLPs. Figure 41(b) illustrates the architecture of the baseline. At training time, MLPs are supervised from the pseudo-labels obtained by graph-based baselines using Mean Squared Error (MSE) score to align the predicted probability distributions to the supervising ones. Missing keysteps and procedural mistakes are predicted from the outputs of the MLP components as in graph-based baselines.
Results
We carried out experiments on the following four scenarios: Install a Wheel, Remove a Wheel, Clean and Lubricate the Chain, First AID - CPR, COVID-19 rapid antigen test. We evaluated our baselines using the calibrated Average Precision (cAP) [30]. Note that, according to this measure, a random baseline would on average achieve a performance of $50\%$.
Table 17 reports the results obtained by our baselines and compares them against those produced by a “uniform” baseline, predicting previous/optional/mistakes/missing/next keysteps with equal probabilities. Results show that the graph-based baseline relying on ground truth annotations significantly outperforms the uniform baseline for some of the tasks, such as previous and missing keystep predictions. This suggests that even simple keystep co-occurrences are informative to some degree of the overall structure of the procedure. The limited performance gains on optional keystep prediction, procedural mistake detection and future keystep prediction highlight the complexity of the task and the need for further research. The end-to-end model trained with instance-level supervision achieves lower or similar performance, trading accuracy for test-time efficiency, due to the absence of an explicit graph. The procedure-level graph-based baseline achieves the best or second-best results for optional keysteps and procedural mistakes, while the end-to-end approach obtains lower performance due to the limited level of supervision it can rely on.
Contributions statement Antonino Furnari led the procedure understanding benchmark, and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Giovanni Maria Farinella contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Luigi Seminara contributed to the annotation guidelines, the baseline design, and paper writing; he also developed tools for data annotation, and the baselines for the benchmark. Francesco Ragusa contributed to the annotation guidelines, the baseline design, and paper writing; he also developed tools for data annotation, and the baselines for the benchmark. Kumar Ashutosh contributed to the annotation guidelines, baseline design, and development of data annotation tools. Michael Wray contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Siddhant Bansal contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Gene Byrne contributed to the task definition, the annotation guidelines and the baseline design. Tushar Nagarajan contributed to the task definition, and the baseline design. Lorenzo Torresani contributed to editing this section.
13.C   Ego-(Exo) Proficiency Estimation
Annotations
The proficiency estimation benchmark consists of two task variants: (1) demonstrator proficiency estimation, where the goal is to estimate the absolute skill level of a participant at the task, and (2) demonstration proficiency estimation, where the goal is to perform fine-grained analysis of a given task execution to identify good actions from the participant and suggest tips for improvement. We now provide a detailed description of the annotations used in each of these task variants.
Demonstrator proficiency estimation. In this task variant, we assign proficiency labels to each person performing activity demonstrations (one label per person). The proficiency labels span four distinct categories: novice, early expert, intermediate expert, or late expert. We find four proficiency classes makes the task challenging but still approachable. Subtle variations between 5 or more levels of proficiency can be insufficiently observable from vision alone, and even difficult for expert annotators to reach consensus. Most levels correspond to experts since Ego-Exo4D videos are dominantly targeted towards expert participants who can perform the task successfully (see Sec. 10).
@@ 13.C   Ego-(Exo) Proficiency Estimation
We derive annotations for this task from participant surveys (see Sec. 11.B) and expert commentary (see Sec. 12.A). Participant surveys contain responses to questions about prior experiences in the task such as “How many years have you been doing this task?", and “Do you have any qualifications/professional training related to the task?" (see Tab. 3 for the complete list). On the other hand, expert commentary is performed by task-specific experts and includes 1 to 10 proficiency scores for each video from the participant (see Sec. 12.A). After consulting with experts hired for each scenario, we designed scenario-specific conversion functions that use the surveys and expert commentaries to produce an estimate of a participant’s proficiency score (see Tab. 19). For example, in basketball and soccer, we use the years of experience to determine skill level since we found this to be an accurate indicator of skill based on analyzing the videos. On the other hand, to determine skill level in bouldering, we use the highest difficulty level of the route solved by the participant.
Note that we exclude the bike repair and health scenarios from the demonstrator proficiency task for different reasons. The distribution of participants for bike repair is heavily skewed towards late expert participants, which would heavily bias the task of demonstrator proficiency. The predominant activity in the health collection is COVID testing, where skill levels are hard to determine due to the simplicity of the task.
See Fig. 42 for a distribution over proficiency scores within each scenario. We split our dataset into train/val/test splits based on the common split shared across benchmarks. The dataset statistics are shown in Tab. 18.
Demonstrator proficiency Demonstration proficiency
Train Val Test Train Val Test
Basketball 575 143 167 146 47 19
Bike repair - - - 41 9 15
Cooking 200 53 86 80 24 39
Dance 380 127 148 80 35 27
Health - - - 42 12 16
Music 138 36 71 94 32 35
Rock Climbing 561 159 230 65 15 22
Soccer 129 43 66 8 3 6
Total 1983 561 768 556 177 179
Demonstration proficiency estimation. For this variant of the task, we leverage temporally localized annotations that include the timestamps of steps demonstrated in the video as well as the proficiency category for each demonstrated step instance (i.e., good execution or a mistake). For this task, we consider all 8 scenarios, as shown in Tab. 18. We derive annotations for this task from expert commentary, where task experts carefully analyze videos and provide timestamped commentary on the participant’s performance (see Sec. 12.A). In particular, given a single timestamped comment from an expert, we annotate whether the comment describes a good execution and/or provides tips for improving the participant’s skill level. See Tab. 20 for example annotations. These annotations are then associated with the timestamp provided with each comment to obtain a list of timestamps for good executions $\{t_{1}^{g},t_{2}^{g},\cdots\}$ and tips for improvement $\{t_{1}^{i},t_{2}^{i},\cdots\}$ in each video. The objective then is to train models that can predict timestamps of good executions or tips for improvement. This amounts to a temporal action localization task for the actions “good executions" and “tips for improvement". While it would be interesting to build systems to further elaborate what these are (e.g., a specific type of tip, or a natural language explanation), we reserve this for future work. Overall, the demonstration proficiency estimation task consists of 556 train / 177 val / 179 test videos (see Tab. 18 for a breakdown per scenarios).
@@
Formal task definition
We now formally define the two problem formulations. In both cases, we are given as input an egocentric video and a set of $M$ exocentric videos of a demonstrator performing a task: $\mathcal{V}=\left(\mathcal{V}_{ego},\mathcal{V}_{exo}^{1-M}\right)$. The ego and exo views are time-aligned.
Demonstrator proficiency estimation. The goal here is to estimate the demonstrator’s skill level from one or more task demonstrations. We wish to learn a function $f$ that can estimate the demonstrator’s proficiency, i.e., $f(\mathcal{V})=y^{prof}\in$novice, early expert, intermediate expert, late expert}.
Demonstration proficiency estimation. On the other hand, the goal in this task variant is to identify parts of the video where the execution was good or needs further improvement. Thus, this time the objective is to learn a function $h$ that can temporally localize instances of good executions and instances of demonstrations that require improvement. Formally, we can express function $h$ as:
$\hat{G},\hat{I}=h(\mathcal{V}),$ (2)
where $\hat{G}=\{t^{g}_{1},t^{g}_{2},\cdots,t^{g}_{|G|}\}$ are the timestamps where the participant shows good task execution, and $\hat{I}=\{t^{i}_{1},t^{i}_{2},\cdots,t^{i}_{|I|}\}$ are the timestamps where the participant needs to improve their skill level.
These tasks inherently benefit from multi-view data. Egocentric video captures fine-grained information about the demonstrator’s hand pose and object interactions, which can be critical to estimating proficiency in tasks that require fine-grained interactions such as cooking (e.g., chopping vegetables) and music (e.g., placement of fingers on the guitar). On the other hand, the exocentric videos provide broader information about the demonstrator’s body pose, which can be highly indicative of proficiency in tasks that require extensive physical motion such as basketball, soccer, and dancing. Our EgoExo4D dataset has $5$ views (1 egocentric view, and $M=4$ exocentric views). We run the proficiency tasks in two settings: one where the exo view is available at test time, and one where it is not. For the latter, benchmarking baseline models with only the egocentric view is important when the target is augmented reality and mobile robotics applications. For the former, benchmarking with both egocentric and exocentric views is informative to capture the multi-view nature of the problem.
Note that the input to the model excludes textual descriptions/narrations of the activity, audio, gaze sensor readings, and any subject information, which would simplify the task significantly at the expense of general applicability since these signals are typically not available for in-the-wild video. We believe the formulation proposed here will encourage the development of video-based methods that (1) do not rely on explicit subject information such as gender, age, ethnicity, etc., and (2) learn to estimate proficiency based on visual cues rather than high-level textual activity descriptions or alternate modalities like sound and eye gaze.
Baselines
Next, we define our proposed baselines for each task in the benchmark.
Demonstrator proficiency estimation. We approach this task as a classification problem with four proficiency classes: novice, early expert, intermediate expert, and late expert. We adopt the TimeSformer [13] model for our experiments. TimeSFormer is a video transformer designed for video action recognition/classification that introduces a novel decoupled spatiotemporal attention mechanism. We train one model on the egocentric view (“ego model"), and a separate model on all 4 exocentric views (“exocentric model"). We resize all videos to 448 pixels along the smallest dimension and use a clip size of 16 frames with a frame rate of 16 FPS. The models are trained to classify individual clips on 8 Quadro RTX 6000 GPUs for 15 epochs. We use the cross-entropy loss as our training objective. At inference time, we employ a late fusion strategy to incorporate information from both egocentric and exocentric video streams. We average the softmax predictions across both egocentric and exocentric models to obtain the final video label prediction. We also average results over three spatial crops during inference following prior work [13]. We report the top-1 accuracy metric.
@@
Demonstration proficiency estimation. We treat this task as a temporal action localization problem with two action categories: ‘good execution’ and ‘needs improvement’. We adopt the ActionFormer [190] model for our experiments. Unlike traditional action localization that defines time windows as outputs, we instead perform timestamp regression since our annotations contain only a single point in time for each good execution or tip for improvement. We accordingly adapt ActionFormer’s post-processing strategy and evaluation metrics. In our task, the predicted timestamps correspond to frames retained after non-maximum suppression (NMS). We remove the regression head of ActionFormer and infer the predicted timestamps from the indices of frames retained after NMS. We also modify the NMS module of ActionFormer to rely on the $L_{1}$-distance between predicted timestamps instead of the tIoU between segments used in [190]. During training, we keep the classification loss from [190] and replace the regression loss with the loss function defined in [75].
We also modify the evaluation setup to use the $L_{1}$-distance between the predicted and ground-truth timestamps instead of tIoU between segments. Therefore, we have metrics thresholded by $L_{1}$-distance in seconds as opposed to tIoU values. We train our models with Omnivore features [45] extracted with a clip size of 32 frames and a stride of 16 frames from all the videos. For the experiments involving multiple views (i.e., multiple exo views or ego + exo views), we simply concatenate the features for all views at each time step.
Accuracy
v1 v2
Ego Exos Ego + Exos Ego Exos Ego + Exos
Method Pretraining Val Test Val Test Val Test Val Test Val Test Val Test
Random - 24.9 24.9 24.9 24.9 24.9 24.9 26.4 26.4 26.4 26.4 26.4 26.4
Majority-class - 31.1 35.3 31.1 35.3 31.1 35.3 32.3 42.4 32.3 42.4 32.3 42.4
TimeSFormer [13] - 42.3 42.3 40.1 48.6 40.8 47.8 40.6 33.9 39.0 47.5 39.9 45.7
TimeSFormer [13] K400 42.9 47.3 39.1 49.4 38.6 48.5 47.2 44.1 37.8 47.0 40.3 46.1
TimeSFormer [13] HowTo100M 46.8 48.5 38.2 38.1 39.7 40.7 45.1 36.7 39.8 46.6 43.7 47.0
TimeSFormer [13] EgoVLP 44.4 48.6 40.6 46.0 39.5 47.2 44.7 43.8 40.5 44.5 39.4 43.5
TimeSFormer [13] EgoVLPv2 45.9 39.5 38.0 46.7 37.8 45.7 46.7 50.4 37.0 47.0 37.1 48.7
Inference with multiple takes per demonstrator
K400 43.3 42.1 38.0 52.4 38.6 52.0 - - - - - -
EgoVLPv2 - - - - - - 48.3 51.0 36.0 47.3 43.1 49.1
Table 21: Demonstrator proficiency estimation benchmark. We report top-1 accuracies for various baselines on the demonstrator proficiency estimation task, on both the v1 and the v2 test set. All models are trained on the v1 training set.
Results
Demonstrator proficiency estimation.
We present results for demonstrator proficiency estimation in Tab. 21. We include two naïve baselines to account for biases in the dataset. The random baseline uniformly samples one skill level at random. The majority-class baseline predicts the majority class within each scenario. While the random baseline performs poorly, the majority-class baseline achieves higher accuracies, which results from scenario-specific biases in the distribution of expertise levels (see Fig. 42). TimeSFormer trained from random initialization on v1 data outperforms the naïve baselines by a significant margin, demonstrating the ability of learned methods to quantify skill levels from videos. In both val and test splits, the ego videos are sufficient to achieve good performance. In the test split, using the exo videos further improves performance in tasks such as bouldering, highlighting the complementary nature of the ego and exo viewpoints. When models are initialized using HowTo100M and EgoVLP pre-training, we observe good improvements using the ego videos, while the performance using exo videos does not improve over random initialization. In contrast, K400 initialization leads to significant improvements across multiple settings on v1 data, and EgoVLPv2 initialization similarly improves performance on v2 data. In almost all cases, we found that fusing the predictions from the ego view and exo views does not lead to improved results, likely due to the simplicity of our late-fusion strategy. In the last rows of Tab. 21, we further report results when evaluating the best models from v1 and v2 respectively on all demonstrations from the same participant (as opposed to a single demonstration). This multi-take analysis benefits the exo model and the ego + exo model fusion on the test split, but fails to improve on other cases, likely due to the simplicity of our late-fusion across multiple takes. Nevertheless, this highlights the potential for developing more complex strategies to study a participant’s skill level across several demonstrations to obtain more accurate skill estimates. We show scenario-specific results in Tab. 22. The TimeSFormer model achieves good performance on v1 data with the egocentric view in cooking since a close-up view of the objects of interest and hand poses is essential to assessing skill in these scenarios. On the other hand, the model performs better on v1 data with the exocentric view in music and basketball since the overall body pose is a useful indicator of proficiency. Unfortunately, the TimeSFormer model fails to improve over the majority-class baseline for soccer scenarios. Evaluating these v1-trained models on the v2 test set underperforms a naive majority-class baseline, highlighting the differences between the data in these versions. Overall, our benchmark presents new challenges for video-based skill understanding and our results highlight the difficulty of the task, suggesting good scope for improvement in future work.
Demonstration proficiency estimation. We present results for the demonstration proficiency estimation task in Tab. 23. We include three naïve baselines along with ActionFormer [190]. The “Random tips/good exec." baseline randomly predicts a tip or a good execution label every $5.97$ seconds, i.e., the average temporal span between adjacent annotations in our dataset. The “Uniform tips" baseline predicts a tip for improvement label every $5.97$ seconds. The “uniform good exec." baseline predicts a good execution label every $5.97$ seconds. We evaluate ActionFormer models trained on ego only, exo only and ego + exo views. All naïve baselines perform poorly on this task. The learned ActionFormer baseline outperforms the naïve baselines by a good margin. However, the absolute mAP scores are fairly low, suggesting that the task is very challenging and has a significant scope for improvement in methods.
Val results
v1 (K400) v2 (EgoVLPv2)
@@ Results
Basketball 36.19 51.43 52.38 55.24 35.66 56.64 56.64 51.75
Cooking 50.00 45.00 35.00 35.00 50.94 56.60 45.28 49.06
Dancing 51.61 55.65 42.74 42.74 43.31 42.52 31.50 32.28
Music 58.97 46.15 69.23 56.41 44.44 69.44 50.00 58.33
Bouldering 0.00 25.31 17.28 17.28 0.00 27.67 18.24 17.61
Soccer 62.50 56.25 75.00 75.00 74.42 65.11 37.21 41.86
Test results
v1 (K400) v2 (EgoVLPv2)
Basketball 38.30 56.74 63.83 62.41 46.71 79.64 71.86 76.65
Cooking 36.11 52.78 33.33 33.33 39.54 19.77 33.72 32.56
Dancing 45.14 46.53 43.06 41.67 46.62 43.24 40.55 42.57
Music 60.47 55.81 76.74 74.42 73.24 56.34 45.07 50.70
Bouldering 19.65 39.30 42.80 42.80 19.57 39.13 43.04 43.04
Soccer 62.50 50.00 41.67 41.67 72.73 65.15 31.82 30.30

### v4

@@ E.4 Procedure Understanding
Keystep assignment (A1) is applied to obtain a pseudo-labeling of the provided video segments when the supervision is at the procedure-level (i.e., segments are unlabeled and only keystep names are provided). This is achieved by means of an EgoVLPv2 model (Lin et al., 2022) pre-trained on ego-exo videos and narrations. Video segments and keystep names are projected to the shared video-language space using EgoVLPv2. We hence assigned each video segment to the closest keystep in the representation space according to the cosine distance. In the problem formulation with instance-level supervision (i.e., when keystep labels are available for all segments during both training and testing), we use ground truth labels instead of those obtained from keystep assignment. Additionally, we provide a baseline where the keystep assignment step is replaced with label predictions from the Keystep Recognition task (A2). Note that in the training set, segments with a confidence score below 20% have been discarded.
The procedural reasoning component (B) creates for each procedure a transition graph based on keystep co-occurrences. In the graph, each node represents a keystep category, while directed edges represent the probability of transitioning from one node to another one. An edge $A\to B$ is assigned the following weight based on statistics collected from the training videos:
@@
End-to-end baseline. This baseline aims to provide an end-to-end approach to perform procedure understanding directly from the input clip. The baseline predicts previous keysteps, optional keysteps, and next keysteps by feeding video segment features extracted with EgoVLPv2 (Pramanick et al., 2023) to three dedicated MLPs. Figure 23(b) illustrates the architecture of the baseline. At training time, MLPs are supervised from the pseudo-labels obtained by graph-based baselines using Mean Squared Error (MSE) score to align the predicted probability distributions to the supervising ones. Missing keysteps and procedural mistakes are predicted from the outputs of the MLP components as in graph-based baselines.
E.5 Proficiency estimation
Annotations for demonstrator proficiency estimation
@@ E.5 Proficiency estimation
We derive annotations for this task from participant surveys (see Section C) and expert commentary (see Section D.1). Participant surveys contain responses to questions about prior experiences in the task such as “How many years have you been doing this task?”, and “Do you have any qualifications/professional training related to the task?” (see Table 16 for the complete list). On the other hand, expert commentary is performed by task-specific experts and includes 1 to 10 proficiency scores for each video from the participant (see Section D.1). After consulting with experts hired for each scenario, we designed scenario-specific conversion functions that use the surveys and expert commentaries to produce an estimate of a participant’s proficiency score (see Table 22). For example, in basketball and soccer, we use the years of experience to determine skill level since we found this to be an accurate indicator of skill based on analyzing the videos. On the other hand, to determine skill level in bouldering, we use the highest difficulty level of the route solved by the participant.
@@
Annotations for demonstration proficiency estimation
Table 23 shows examples of expert comments and corresponding annotation tags derived from them indicating whether the comment suggests a good execution and/or tips for improvements.
@@
Baseline implementation details
Demonstration proficiency estimation: We use the TimeSFormer (Bertasius et al., 2021) architecture for the baseline. TimeSFormer is a video transformer designed for video action recognition/classification that introduces a novel decoupled spatiotemporal attention mechanism. We resize all videos to 448 pixels along the smallest dimension and use a clip size of 16 frames with a frame rate of 16 FPS. The models are trained to classify individual clips using the cross-entropy loss on 8 Quadro RTX 6000 GPUs for 15 epochs.
@@
Demonstration proficiency estimation: We adopt ActionFormer (Zhang et al., 2022), a video action localization model for our experiments. Unlike traditional action localization that defines time windows as outputs, we instead perform timestamp regression since our annotations contain only a single point in time for each good execution or tip for improvement. We accordingly adapt ActionFormer’s post-processing strategy and evaluation metrics. In our task, the predicted timestamps correspond to frames retained after non-maximum suppression (NMS). We remove the regression head of ActionFormer and infer the predicted timestamps from the indices of frames retained after NMS. We also modify the NMS module of ActionFormer to rely on the $\mathcal{L}_{1}$-distance between predicted timestamps instead of the tIoU between segments used in (Zhang et al., 2022). During training, we keep the classification loss from (Zhang et al., 2022) and replace the regression loss with the loss function defined in (Kwak et al., 2020). We train our models with Omnivore features (Girdhar et al., 2022) extracted with a clip size of 32 frames and a stride of 16 frames from all the videos.
Scenario-specific results on demonstrator proficiency estimation
We show scenario-specific results in Table 24. TimeSFormer achieves good performance with the egocentric view in cooking since a close-up view of the objects of interest and hand poses is essential to assessing skill in these scenarios. On the other hand, the model performs better with the exocentric view in bouldering on the test data since the overall body pose is a useful indicator of proficiency. Unfortunately, it fails to improve over the majority-class baseline is most scenarios on the test splits except basketball, highlighting a distribution shift between the val and test splits.
Val Test
@@
Basketball 35.66 56.64 56.64 51.75 46.71 79.64 71.86 76.65
Cooking 50.94 56.60 45.28 49.06 39.54 19.77 33.72 32.56
Dancing 43.31 42.52 31.50 32.28 46.62 43.24 40.55 42.57
Music 44.44 69.44 50.00 58.33 73.24 56.34 45.07 50.70
Bouldering 0.00 27.67 18.24 17.61 19.57 39.13 43.04 43.04
Soccer 74.42 65.11 37.21 41.86 72.73 65.15 31.82 30.30